SEMICLASSICAL ANALYSIS OF A NONLINEAR EIGENVALUE PROBLEM AND NON ANALYTIC HYPOELLIPTICITY

BERNARD HELFFER DIDIER ROBERT XUE PING WANG

Abstract. We give a semiclassical analysis of a nonlinear eigenvalue problem arising from the study of the failure of analytic hypoellipticity and obtain a general family of hypoelliptic, but not analytic hypoelliptic operators.